A 83 kg adult sits at one end of a 7.0 m long board. His 30 kg child sits on the other end.

a.)Where should the pivot be placed so that the board is balanced, ignoring the board's mass?

b.Find the pivot point if the board is uniform and has a mass of 12 kg
Answer #1

A)

Taking moment about point F,

83 x = 30 (7 - x)

83x = 210 - 30x

X = 1.86 m

B)

83x = 12(3.5 - x) + 30(3.5 + 3.5 - x)

83x = 42 - x + 210 - 30x

X = 2.21 m

Comment in case any doubt please rate my answer...